# PsyMem: Fine-grained psychological alignment and Explicit Memory Control for Advanced Role-Playing LLMs

### v1

@@ 1 Introduction
Recent advancements in large language models (LLMs) (Achiam et al., 2023; Anthropic, 2024a; Anil et al., 2023) have unlocked new possibilities for implementing role-playing systems through language-based behavior simulation (Shanahan et al., 2023). Role-playing systems demonstrate societal value across applications like interactive gaming (Wang et al., 2023a) and AI counseling (Zhang et al., 2024), particularly in enabling social simulation. Such simulation allows modeling of human interactions and societal dynamics (Park et al., 2023), offering insights into policy design and conflict mediation (Zhou* et al., 2024). To achieve trustworthy social simulation, LLM-based role-playing must preserve persistent traits, memories, and behavioral logic to ensure agents reliably mirror specified attributes in evolving social contexts.
@@
In this work, we propose a widely recognized modern psychological evaluation framework (Costa and McCrae, 2008; Schwartz, 1992; Thomas, 2008; Zimbardo and Ruch, 1975) and a graph-structured memory mechanism to construct the PsyMem framework, enabling LLMs to achieve exceptional role-playing abilities. Drawing from modern psychology and neuroscience, we enhance the description of character attributes by leveraging both latent psychological attributes and explicit behavioral patterns (Snyder, 1983; Kahneman, 2011), with personality and values corresponding to latent psychological attributes, and behavioral decision-making, social interaction, and leadership representing explicit behavioral patterns. Overall, we describe characters using 26 quantitative indicators and a small amount of textual descriptions. We extract real data from novels and construct a publicly available role-playing dataset. As shown in Table 1, this dataset contains 5,414 established characters, 38,962 conversations, and 536,636 utterances, much larger than existing role-playing datasets.
@@
  • Beyond textual descriptions, we present a comprehensive set of attributes that bridges intrinsic (e.g., values, emotions) and extrinsic (e.g., behaviors, social interactions) character dimensions, operationalized via 26 measurable indicators. This attribute set captures both the internal motivations and external behavioral manifestations of characters.
  • We introduce a memory alignment training paradigm that establishes explicit cognitive anchoring between role-playing responses and character memory. This methodology enhances role-playing through a dual-control mechanism, enabling precise regulation of generative outputs via both trait-based attributes (e.g., personality dimensions) and episodic memory constraints.
  • We introduce the use of role-based general supervised fine-tuning data (transforming general SFT data into role-playing style SFT data), demonstrating that incorporating such data effectively enhances the performance of Role-Playing LLMs.
@@ 2.1 LLMs-based Role-Playing
Our work belongs to parametric Training. We argue that a character’s behavior is jointly influenced by their attributes, memory, and the scenario. Drawing from modern psychology, we segment character attributes using contemporary psychological frameworks and construct a dataset containing 38,962 conversations (536,636 utterances) from real novels, leveraging both character attributes and memory to enhance the model’s role-playing ability.
@@
Existing LLMs-based Role-Playing commonly relies on the internal knowledge within the model weights to implicitly mimic character memory (Zhou et al., 2023; Lu et al., 2024; Li et al., 2023). For example, Shao et al. (2023) and others push the boundaries of knowledge to protect character memory, while Lu et al. (2024) and others use a self-alignment approach to distinguish the different memory backgrounds of various characters. Relying on the internal knowledge within the model is effective for famous characters but fails for infamous and unseen characters. Some works apply Retrieval-Augmented Generation (RAG) (Ram et al., 2023) which dynamically retrieving external contexts to explicitly mimic character’s memory. For example, Zhong et al. (2024) adopts a memory update mechanism inspired by Ebbinghaus’s forgetting curve theory, emphasizing the recall of relevant memories to understand the user’s personality. However, as shown in Table 2, existing role-playing models tend to perform better when simulating famous characters, while their performance significantly drops for ordinary characters that rely on external memory.
@@
The current evaluation process consists of dialogue generation and dialogue assessment. Due to its efficiency and completeness, automated dialogue generation is widely adopted. Evaluation methods are generally categorized into three types: Metric-based Evaluation(Li et al., 2023; Wang et al., 2023b), Human Evaluation(Zhou et al., 2023), and “LLMs as Judges”(Shao et al., 2023; Lu et al., 2024). Metric-based evaluation primarily measures the model’s alignment with standard responses, such as original text from novels or manually annotated content. While human evaluation is more precise, its high cost and lack of reproducibility limit its scalability. In contrast, “LLMs as Judges”, with its efficiency, low cost, and strong scalability, is increasingly becoming the preferred choice.
@@ 3 Dataset Architecture
An increasing number of role-playing datasets (Ran et al., 2024; Cui et al., 2023) have begun incorporating modern psychological indicators into their characterization frameworks. However, previous studies typically introduced only a limited selection of psychological attributes as supplementary descriptors, without comprehensively integrating psychological systems into character portrayals (Yu et al., 2024; Wang et al., 2024). Moreover, existing datasets predominantly focus on character attributes, overlooking the significant influence memory has on characters’ behaviors, thought processes, and speech styles(Stumpf and Dunbar, 1991; Ravlin and Meglino, 1987; Pickering and Garrod, 2004). This limitation becomes particularly evident when role-playing characters possess limited background knowledge during the pretraining stage (see Figure 1).
@@
where $N$ is the number of data sample. Each sample is a dialogue turn where $R_{i}$ is the role for response; $P_{i}$ is the profile of $R_{i}$, featuring 26 quantitative psychological dimensions; $C_{i}$ is the previous dialogue turns before the current dialogue. $x_{i}$ is the query to $R_{i}$ and $y_{i}$ is the corresponding response. $M_{i}$ is the retrieved memory response of role $R_{i}$ with respect to query $x_{i}$ and the previous dialogue $C_{i}$.
In this work, the dataset $D_{\mathrm{RP}}$ contains 38,962 samples, extracted from 539 novels characterized by a wide range of genres, rich content, and narrative perspectives of third person.(The genre distribution is presented in Figure 2.)
@@ 3.1 Quantifiable Character Attributes
Explicit behavioral patterns manifest externally in observable actions, interactions, and communication styles. We capture these explicit patterns using:
@@ 3.2 Character Memory Integration
Therefore, we introduce character memory data $M$ to the role playing dataset to enable the memory simulation capability of role playing model. Specifically, we first build a knowledge graph based on the novel via GraphRAG (Edge et al., 2024), and then retrieve relevant memory for the current dialogue via entity mentions, relational triples, and raw dialogue chunks associated with specific scenes. Memory alignment is evaluated on two key dimensions: correctness and rationality. Correctness assesses whether the model accurately utilizes stored character information in response to queries, while rationality examines logical coherence and contextual appropriateness. Additionally, simple summaries or vague memory references are identified as non-rational uses of memory. To further enhance robustness, irrelevant memory items are deliberately introduced as noise, promoting stable, realistic, and consistent character portrayals reflective of human memory dynamics.
@@
3.3 Role-based SFT Data
We propose a Role-based general supervised fine-tuning (SFT) data construction approach. Dialogue samples from the Pure-Dove (PD) dataset (Daniele and Suphavadeeprasit, 2023) are grouped randomly, and LLMs infer each speaker’s latent psychological attributes and explicit behavioral patterns. Scene-specific information is dynamically generated using fine-tuned LLMs based on dialogue contexts, significantly improving character consistency and alignment during interactions, as demonstrated by our experimental evaluations.
Similar to Equation 1, we construct a dataset defined as:
@@ 3.3 Role-based SFT Data
Where $R_{i}$ denotes the role responsible for the response, and $P_{i}$ is the profile of $R_{i}$ defined by 26 quantitative psychological dimensions. $C_{i}$ represents the preceding dialogue context, $x_{i}$ is the current query directed to $R_{i}$, and $y_{i}$ is the corresponding response. These $(x_{i},C_{i},y_{i})$ triples are drawn from general supervised fine-tuning datasets. In addition, $M_{i}$ refers to the retrieved memory response relevant to role $R_{i}$ with respect to $x_{i}$ and $C_{i}$.
@@ 4.1 Training
In this work, we transfer general LLMs to role-playing models via supervised fine-tuning. Role-playing presents a unique challenge for language models: maintaining both general capabilities and character-specific attributes simultaneously. This challenge exemplifies the fundamental tension in transfer learning between preserving general knowledge and specializing for domain-specific tasks. Traditional fine-tuning approaches (Kotha et al., 2024) often suffer from catastrophic forgetting, where optimization for character fidelity compromises the model’s general capabilities. Inspired by the Dual-stage Mixed Fine-tuning (DMT) (Dong et al., 2023) method, we introduce a two-stage training strategy for role-playing, effectively balance character fidelity and the preservation of general language understanding capabilities, as illustrated in Figure 3b.
Specifically, we introduce a two-stage progressive training framework to systematically develop LLMs into specialized Role-Playing Models: Stage 1 establishes core role-playing capabilities through fine-tuning on data emphasizing fundamental role characteristics. Stage 2 enhances specialization via memory-augmented fine-tuning, incorporating long-term memory contexts and general role-based information. This approach ensures a balanced development of character-specific attributes while simultaneously preserving the model’s general capabilities.
Stage 1: Foundational Role-Playing Capacity Development.
In this stage, we train LLMs to learn foundational role-playing ability based only on quantifiable attributes and basic information. We fine-tune the model using a subset $D_{\mathrm{RP1}}$ sampled from the overall role-playing dataset to establish its fundamental role-playing abilities. The objective function is formally defined as:
@@
where ()($\cdot$)( ⋅ ) in $D_{\mathrm{RP}}$ represents the tuple $(R,P,C,x,y)$, $D_{\mathrm{RP1}}$ is a subset sampled from the overall role-playing dataset $D_{\mathrm{RP}}$, excluding the retrieved memory $M$.
Stage 2: Integrated Memory-Augmented Specialization. To achieve integrated memory-augmented specialization, we further specialize the model by fine-tuning it on a distinct subset from the role-playing dataset $D_{\mathrm{RP}}$, augmented with long-term memory contexts retrieved via graph-based methods, and combined with general role-based SFT data $D_{\mathrm{RB}}$. This stage employs a weighted combination of two loss terms, explicitly designed to mitigate forgetting of general capabilities while simultaneously enhancing the model’s ability to effectively leverage long-term memory contexts:
[…]
@@
[…]
where ()($\cdot$)( ⋅ ) in $D_{\mathrm{RP2}}$ represents the tuple $(R,P,M,C,x,y)$ and in $D_{\mathrm{RB}}$ represents $(R,P,C,x,y)$. $M$ denotes the retrieved memory response for role $R$ regarding the query $x$ and prior dialogue context $C$, obtained via graph-based retrieval methods and subsequently filtered by LLMs.
@@ 4.2 Evaluation
We evaluate role-playing capabilities of LLMs by adopting the widely used “LLMs as Judges” approach (Shao et al., 2023; Lu et al., 2024), as illustrated in Figure 3c. The evaluation process is structured into two main categories: Character-independent Capabilities and Character Fidelity. Additionally, in the Tabel 4, we further discuss the discrepancies between LLM-based judgments and human evaluations across 26 quantitative indicators to validate the effectiveness of the “LLMs as Judges” approach.
4.2.1 Character-independent Capabilities
Following previous research(Chen et al., 2024; Yu et al., 2024), we measure Character-independent Capabilities using three metrics: Human-likeness, Consistency and Coherence. In the Human-likeness task, we use GPT-4o with few-shot examples to determine whether a dialogue was generated by a human, ultimately computing the accuracy score. For consistency, we evaluate the coherence between different parts of the conversation to assess whether the role-playing remains consistent throughout long dialogues. Additionally, for coherence, we assess how logically connected and contextually appropriate the responses are across the entire dialogue, ensuring the conversation flows naturally.
4.2.2 Character Fidelity
Character Fidelity assesses how accurately an LLM portrays specific role-playing characters. As described in Section 3.1, we employ modern psychological quantification frameworks defined in our dataset design, covering four primary dimensions: Personality (Big Five Personality Model), Values (Schwartz’s Theory of Basic Values), Social and Leadership Styles (Zimbardo’s psychological principles), and Behavioral Decision-making (Thomas-Kilmann Conflict Mode Instrument, TKI). These dimensions are quantified into numeric scales using standardized psychological evaluation methods. Additionally, we separately evaluate the model’s memory, focusing on the correctness and rationality of memory alignment. The Memory dimension is measured using accuracy scores derived from a true/false evaluation format.
To calculate similarity between role-playing characters, we employ Normalized Euclidean Distance (NED), computed as follows:
@@ 4.2.2 Character Fidelity
where $\Delta X=X_{\max}-X_{\min}$, $n$ is the number of subdimensions, $X_{\mathrm{ref}}$ is the standardized reference value, and $X_{i}$ represents the measurement for the $i$-th subdimension, with possible values from $Z=\{1,2,3,4,5\}$. This formulation is robust, as it normalizes the mean squared deviation across different scales, yielding a more interpretable similarity metric.

### v2

@@ 1 Introduction
Recent advancements in large language models (LLMs) (Achiam et al., 2023; Anthropic, 2024a; Anil et al., 2023) have unlocked new possibilities for implementing role-playing systems through language-based behavior simulation (Shanahan et al., 2023). Role-playing systems demonstrate societal value across applications like interactive gaming (Wang et al., 2023a) and AI counseling (Zhang et al., 2024), with particularly potential in enabling high-reliable social simulation. Such simulation allows modeling of human interactions and societal dynamics (Park et al., 2023), offering insights into evidence-based policy design and conflict mediation (Zhou* et al., 2024). To achieve trustworthy social simulation, LLM-based role-playing must preserve persistent traits, memories, and behavioral logic to ensure agents reliably mirror specified attributes in evolving social contexts.
@@
In this work, we propose a widely recognized modern psychological evaluation framework (Costa and McCrae, 2008; Schwartz, 1992; Thomas, 2008; Zimbardo and Ruch, 1975) and a graph-structured memory mechanism to construct the PsyMem framework, enabling LLMs to achieve exceptional role-playing abilities. Drawing from modern psychology and neuroscience, we enhance the description of character attributes by leveraging both latent psychological attributes and explicit behavioral patterns (Snyder, 1983; Kahneman, 2011), with personality and values corresponding to latent psychological attributes, and behavioral decision, social interaction, and leadership representing explicit behavioral patterns. Overall, we extract real data from novels and construct a large scale role-playing dataset, using 26 quantitative indicators and a small amount of textual descriptions to describe characters. As shown in Table 1, this dataset contains 5,414 established characters, 38,962 conversations, and 536,636 utterances, much larger than existing role-playing datasets.
@@
  • Beyond textual descriptions, we present a comprehensive attribute set that bridges intrinsic (e.g., values, emotions) and extrinsic (e.g., behaviors, social interactions) character dimensions, operationalized via 26 measurable indicators. This attribute set captures both the internal motivations and external behavioral manifestations of characters.
  • We introduce a memory alignment training paradigm that establishes explicit cognitive anchoring between role-playing responses and character memory. This methodology enhances role-playing through a dual-control mechanism, enabling precise regulation of generative outputs via both attributes (e.g., personality dimensions) and episodic memory constraints.
  • We introduce the use of role-playing style general supervised fine-tuning data (transforming general SFT data into role-playing style), demonstrating that incorporating such data effectively enhances the performance of Role-Playing LLMs.
@@ 2.1 LLMs-based Role-Playing
Our work belongs to parametric training. We argue that a character’s behavior is jointly influenced by their attributes, memory, and the scenario. Drawing from modern psychology, we segment character attributes using contemporary psychological frameworks and construct a dataset containing 38,962 conversations (536,636 utterances) from real novels, leveraging both character attributes and memory to enhance role-playing ability.
@@
Existing LLMs-based Role-Playing commonly relies on the internal knowledge within the model weights to implicitly mimic character memory (Zhou et al., 2023; Lu et al., 2024; Li et al., 2023). For example, Shao et al. (2023) and others Lu et al. (2024); Yu et al. (2024) push the boundaries of knowledge to protect character memory, while Lu et al. (2024) and others use a self-alignment approach to distinguish the different memory backgrounds of various characters. Relying on the internal knowledge within the model is effective for famous characters but fails for unfamous and unseen characters. Some works apply Retrieval-Augmented Generation (RAG) (Ram et al., 2023) to dynamically retrieve external contexts for explicitly mimicking character memory, such as Zhong et al. (2024) who adopts memory update mechanisms based on Ebbinghaus’s forgetting curve theory. However, as shown in Table 2, existing role-playing models tend to perform better when simulating famous characters, while their performance significantly drops for ordinary characters that rely on external memory.
@@
The current evaluation process consists of dialogue generation and dialogue assessment. Due to its efficiency and completeness, automated dialogue generation is widely adopted. Evaluation methods are generally categorized into three types: Metric-based Evaluation(Li et al., 2023; Wang et al., 2023b), Human Evaluation(Zhou et al., 2023), and “LLMs as Judges”(Shao et al., 2023; Lu et al., 2024). Metric-based evaluation primarily measures the model’s alignment with standard responses, such as original text from novels or manually annotated content. While human evaluation is more precise, its high cost and lack of reproducibility limit its scalability. In contrast, “LLMs as Judges”, with its efficiency, low cost, and strong scalability, is increasingly becoming the preferred choice. The experiments in Appendix (see Table 7) demonstrate the effectiveness of the “LLMs as Judges” paradigm for role-playing evaluation.
@@ 3 Dataset Architecture
An increasing number of role-playing datasets (Ran et al., 2024; Cui et al., 2023) have begun incorporating modern psychological indicators into their characterization frameworks. However, previous studies typically introduced only a limited selection of psychological attributes as supplementary descriptors, without comprehensively integrating psychological systems into character portrayals (Yu et al., 2024; Wang et al., 2024). Moreover, existing datasets predominantly focus on character attributes, overlooking the significant influence memory has on characters’ behaviors, thought processes, and speech styles(Stumpf and Dunbar, 1991; Ravlin and Meglino, 1987; Pickering and Garrod, 2004). This limitation becomes particularly evident when LLMs can hardly access the background knowledge of the role-playing characters during the pretraining stage (see Figure 1).
@@
where $N$ is the number of data sample. Each sample is a real dialogue turn collected from novels. $R_{i}$ is the role for response; $P_{i}$ is the profile of $R_{i}$, including basic information and 26 quantitative psychological dimensions; $C_{i}$ consists of the dialogue turns preceding the current turn, as well as the scene information. $x_{i}$ is the query to $R_{i}$ and $y_{i}$ is the corresponding response. $M_{i}$ denotes the memory retrieved for role $R_{i}$, conditioned on the query $x_{i}$ and the context $C_{i}$, which comprises the preceding dialogue turns and the scene.
In this work, the dataset $\text{D}_{\text{RP}}$ contains 38,962 samples, extracted from 539 novels characterized by a wide range of genres, rich content, and narrative perspectives of third person.(The genre distribution is presented in Figure 2.)
@@ 3.1 Quantifiable Character Attributes
Explicit behavioral patterns manifest in observable actions, interactions, and communication styles. We capture these explicit patterns using:
@@ 3.2 Character Memory Integration
Therefore, we introduce character memory data $M$ to the role playing dataset to enable the memory simulation capability of role playing model. Specifically, we first build a knowledge graph for each novel via GraphRAG (Edge et al., 2024), and then retrieve relevant memory context for the current dialogue in local mode (see Appendix B.1 for details). Memory alignment is evaluated on two key dimensions: correctness and rationality. Correctness assesses whether the model accurately utilizes stored character information in response to queries, while rationality examines logical coherence and contextual appropriateness. Additionally, simple summaries or vague memory references are identified as non-rational uses of memory. To further enhance robustness, irrelevant memory items are deliberately introduced as noise, promoting stable, realistic, and consistent character portrayals reflective of human memory dynamics.
@@
3.3 Synthetic Role-playing Data
In addition to $\text{D}_{\text{RP}}$, we propose a synthetic role-playing data construction approach that transforms general supervised fine-tuning (SFT) data into role-playing style dialogues, resulting in the dataset $\text{D}_{\text{RP}}^{\text{synth}}$. The construction pipeline consists of three steps: 1) We randomly sample multi-turn dialogues from the Pure-Dove (PD) dataset (Daniele and Suphavadeeprasit, 2023) and sequentially group every 100 consecutive dialogue turns to form speaker-specific clusters. Each cluster containing approximately 100 dialogue turns and is then assigned to represent a distinct synthetic speaker. 2) For each synthetic speaker, we use GPT-4o to analyze the 100 dialogue turns and annotate the speaker’s psychological profile according to our 26 quantitative dimensions. 3) Based on the dialogue context, we employ GPT-4o to generate appropriate scene-specific information to enhance contextual coherence. The resulting dataset $\text{D}_{\text{RP}}^{\text{synth}}$ is formulated as:
@@ 3.3 Synthetic Role-playing Data
where $R_{i}$ denotes the role responsible for the response, and $P_{i}$ is the profile of $R_{i}$ defined by basic information and 26 quantitative psychological dimensions. $C_{i}$ represents the preceding turns and scene information. $x_{i}$ and $y_{i}$ are the current query, and the corresponding response drawn from the Pure-Dove (PD) datasets, respectively.
By incorporating role-playing style general SFT data $\text{D}_{\text{RP}}^{\text{synth}}$, the proposed PsyMem framework enables the model to develop stronger role-playing capabilities, as Figure 4 shows.
@@ 4.1 Training
In this work, we transfer general LLMs to role-playing models via supervised fine-tuning. Role-playing presents a unique challenge for language models: maintaining both general capabilities and character-specific attributes simultaneously. This challenge exemplifies the fundamental tension in transfer learning between preserving general knowledge and specializing for domain-specific tasks. Traditional fine-tuning approaches (Kotha et al., 2024) often encounter a core multi-task conflict during optimization. In the context of role-playing tasks, this conflict arises from the attempt to optimize two fundamentally distinct objectives—namely, character profile alignment and memory adherence—within a unified framework. Inspired by the Dual-stage Mixed Fine-tuning (DMT) (Dong et al., 2023) method, we introduce a two-stage training strategy for role-playing, effectively balancing the capabilities of character profile alignment and memory adherence.
Specifically, stage 1 establishes core role-playing capabilities by training the model on character profile-based dialogues, enabling the model to learn fundamental character attribute alignment. Stage 2 further trains the model on memory-augmented role-playing dataset $\text{D}_{\text{RP}}$ and role-playing style general fine-tuning dataset $\text{D}_{\text{RP}}^{\text{synth}}$ to improve role-playing precision, as illustrated in Figure 3(b). As demonstrated in Section 5.3, our two-stage training strategy significantly outperforms vanilla single-stage training for role-playing performance, validating the necessity of this progressive specialization strategy.
Stage 1: Foundational Role-Playing Capacity Development
In this stage,we train LLMs to learn foundational role-playing abilities based on quantifiable character attributes and dialogue contexts without memory augmentation. We fine-tune the model using a subset $\text{D}_{\text{RP1}}$ sampled from the overall role-playing dataset $\text{D}_{\text{RP}}$ to establish its fundamental role-playing abilities. The objective function is formally defined as:
@@
where $(\cdot)$ in $\text{D}_{\text{RP1}}$ denotes the tuple $(R,P,C,x,y)$ and $\text{D}_{\text{RP1}}$ excludes the retrieved memory component $M$ from the original dataset $\text{D}_{\text{RP}}$.
Stage 2: Memory-Augmented Role-Playing Specialization. In this stage, we enhance the model’s role-playing capabilities through two training components: 1) memory-augmented role-playing data $\text{D}_{\text{RP}}$, which incorporates long-term memory contexts retrieved via graph-based methods to enable explicit memory alignment, and (2) synthetic role-playing style data $\text{D}_{\text{RP}}^{\text{synth}}$, which improves general role-playing abilities. This stage employs a weighted combination of two corresponding loss terms to simultaneously enhance memory-controlled response generation while preserving the model’s foundational capabilities:
$\theta^{*}_{2}=\arg\min_{\theta}\left\{\alpha\mathcal{L}_{\mathrm{RP}}(\theta)+\mathcal{L}_{\mathrm{RP}}^{\mathrm{synth}}(\theta)\right\}$
@@
$\mathcal{L}_{\mathrm{RP}}^{\mathrm{synth}}(\theta)=\mathbb{E}_{(\cdot)\sim\text{D}_{\text{RP}}^{\text{synth}}}\left[\mathcal{L}\left(f_{\theta}(R,P,C,x),y\right)\right]$
where $(\cdot)$ in $\text{D}_{\text{RP}}^{\text{synth}}$ represents $(R,P,C,x,y)$ and $(\cdot)$ in $\text{D}_{\text{RP}}$ represents the tuple $(R,P,M,C,x,y)$. $M$ denotes the retrieved memory response for role $R$ regarding the query $x$ and the context $C$ (which includes the preceding dialogue turns and scene information), obtained via graph-based retrieval methods and subsequently filtered by LLMs.
@@ 4.2 Evaluation
We evaluate role-playing capabilities of LLMs by adopting the widely used “LLMs as Judges” approach***We validate the reliability of our LLM-based evaluation framework by comparing human-evaluation and GPT-4o judgments in Appendix (see Table 7) (Shao et al., 2023; Lu et al., 2024), as illustrated in Figure 3(c). The evaluation process is structured into two main categories: Character-independent Capabilities and Character Fidelity.
Character-independent Capabilities.
Following previous research (Chen et al., 2024; Yu et al., 2024) , we measure Character-independent Capabilities using three metrics: Human-likeness, Consistency and Coherence. In the Human-likeness task, we use GPT-4o with few-shot examples to determine whether a dialogue was generated by a human, ultimately computing the accuracy score. For consistency, we evaluate the coherence between different parts of the conversation to assess whether the role-playing remains consistent throughout long dialogues. For coherence, we assess how logically connected and contextually appropriate the responses are across the entire dialogue, ensuring the conversation flows naturally.
Character Fidelity.
Character Fidelity assesses how accurately an LLM portrays specific role-playing characters. As described in Section 3.1, we employ modern psychological quantification frameworks defined in our dataset design, covering five primary dimensions: Personality (Big Five Personality Model), Values (Schwartz’s Theory of Basic Values), Social and Leadership Styles (Zimbardo’s psychological principles), Behavioral Decision (Thomas-Kilmann Conflict Mode Instrument, TKI), and Memory. The former four dimensions are quantified into numeric scales ranging from 1 to 5 using standardized psychological evaluation methods. The Memory dimension, distinct from the psychological attributes, is evaluated using accuracy scores from true/false assessments.
Evaluation Process.
The evaluation process commences with models generating 15-turn dialogues per scenario (see Section 5.1 for scenario detail), embodying the specified character profile. For character-independent capabilities evaluation, we utilize these generated dialogues: 1) Human-likeness and Coherence are assessed by GPT-4o (true/false based on rules), yielding a positive assessment rate (e.g., the percentage deemed human-like or coherent). 2) Consistency is measured by the score difference in the four psychological dimensions between dialogue halves, yielding a stability score (e.g., where a smaller difference yields a higher consistency score, scaled from 0 to 1). For character fidelity evaluation, GPT-4o serves as the judge, assessing each 15-turn dialogue. Performance is scored across five dimensions: 1) The first four dimensions (Personality, Values, etc.,) are evaluated by GPT-4o on dialogues, using 1-NED (introduced in the following paragraph) for performance; 2) For the fifth dimension, memory alignment, we use GPT-4o to judge response correctness (true/false) against memory context and reference answer, yielding an accuracy score (i.e., the proportion of correctly aligned responses).
Metrics.
We quantify role-playing fidelity for Personality, Values, Social & Leadership, and Behavioral Decision using the 1-NED metric. Derived from Normalized Euclidean Distance (NED), a higher 1-NED score indicates greater similarity to the target profile. The 1-NED score is computed as:
@@
where $\Delta\!X=\!\!X_{\max}\!\!-\!X_{\min}$, with $X_{\max}$ and $X_{\min}$ being the maximum (5) and minimum (1) values of the evaluation scale, respectively. A score of 0, used to denote uncertainty, is excluded from this range. $n$ is the number of subdimensions, $X_{i}$ represents the measurement for the $i$-th subdimension, and $X_{\mathrm{ref}}$ is the corresponding value contained in role profile $P_{i}$ (see Eq.(1) and Eq.(2)), with possible values from $\{1,2,3,4,5\}$. This formulation is robust, as it normalizes the mean squared deviation across different scales, yielding a more interpretable similarity metric.